OpenStudy (anonymous):

Find the fourth term in the expansion of (3x-y^2)^6.

OpenStudy (anonymous):

The fourth term is 20a^3 b^3. Substituting in a=3x and b=-y^2, this is 20(3x)^3 (-y^2)^3 = -20*3^3*x^3*y^6 = -540 x^3 y^6

OpenStudy (anonymous):

im confused on what my answer will be. @wmckinely

OpenStudy (anonymous):

nvm i got it..